Ariel:

Sara left a bin outside in her garden to collect rainwater. She notices that 1 over 8 gallon of water fills 2 over 3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin. Show your work. Explain your answer in words.

Narad:

Let the volume of the bin (which is the volume of water) \[=x \] gallons The equation is \[2/3x = 1/8\] (two thirds of the bin is equivalent to one eigth gallon) Multiply both sides of the equation by\[3/2\] \[3/2* 2/3x = 1/8 *3/2\] \[x=3/16\] gallon The volume of the bin is three sixteenth of a gallon